On January 1, 2020, Pharoah Company had the following stockholders equity accounts. Common Stock ($26 par value, 58,500 sharJournalize the transactions and the closing entries for net income and dividends. (Credit account titles are automatically inEnter the beginning balances, and post the entries to the stockholders equity accounts. (Use T-accounts.) (Note: Open additiPrepare a stockholders equity section at December 31. (Enter account name only and do not provide descriptive information.)
Answer #1
Pharoah Corporation                              -  
Date Account Debit Credit Calculation
1-Feb Cash Dividends            58,500 =58,500*1
Dividends Payable            58,500
1-Mar Dividends Payable            58,500
Cash            58,500
1-Apr No Entry
In Stock split, only par value per share changes to $13 per share (half of $26 par)
1-Jul Stock Dividends          245,700 =17,550*14
Common Stock Dividend Distributable          228,150 =17,550*13
Paid-in Capital in Excess of Par—Common Stock            17,550
31-Jul Common Stock Dividend Distributable          228,150
Common Stock          228,150
1-Dec Cash Dividends            53,820 Below wkgs
Dividends Payable            53,820
31-Dec Income Summary          359,000
Retained Earnings          359,000
31-Dec Retained Earnings          245,700
Stock Dividends          245,700
31-Dec Retained Earnings          112,320 =58,500+53,820
Cash Dividends          112,320
Workings
Cash Dividends
Opening Shares            58,500
Stock Split - 2 for 1            58,500
Subtotal          117,000
15% Stock Dividends            17,550 =117,000*15%
Number of Common Shares Issued          134,550
Cash Dividend per share                 0.40
Cash Dividend Declared            53,820 =134,550*0.40

Pharoah Corporation T Accounts Date Debit Credit Common Stock $1,521,000 31-Jul $ 228,150 Op bal Balance Date Debit Credit BaPharoah Corporation Partial Balance Sheet As of December 31, 2020 1,749,150 Stockholders Equity Paid in Capital Capital Stock